# LoRA Users Beware: A Few Spurious Tokens Can Manipulate Your Finetuned Model

### v1

@@ Abstract
Parameter Efficient FineTuning (PEFT), such as Low-Rank Adaptation (LoRA), aligns pre-trained Large Language Models (LLMs) to particular downstream tasks in a resource-efficient manner. Because efficiency has been the main metric of progress, very little attention has been put in understanding possible catastrophic failures. We uncover one such failure: PEFT encourages a model to search for shortcut solutions to solve its fine-tuning tasks. When very small amount of tokens, e.g., one token per prompt, are correlated with downstream task classes, PEFT makes any pretrained model rely predominantly on that token for decision making. While such spurious tokens may emerge accidentally from incorrect data cleaning, it also opens opportunities for malevolent parties to control a model’s behavior from Seamless Spurious Token Injection (SSTI). In SSTI, a small amount of tokens correlated with downstream classes are injected by the dataset creators. At test time, the finetuned LLM’s behavior can be controlled solely by injecting those few tokens. We apply SSTI across models from three families (Snowflake Arctic, Apple OpenELM, and Meta LLaMA-3) and four diverse datasets (IMDB, Financial Classification, CommonSense QA, and Bias in Bios). Our findings reveal three astonishing behaviors. First, as few as a single token of SSTI is sufficient to steer a model’s decision making. Second, for light SSTI, the reliance on spurious tokens is proportional to the LoRA rank. Lastly, with aggressive SSTI, larger LoRA rank values become preferable to small rank values as it makes the model attend to non-spurious tokens, hence improving robustness.
@@ 1 Introduction
Large language models (LLMs) have achieved impressive performance across a range of natural language processing tasks. However, their generalization can at times be fragile, particularly when training data contains spurious correlations—patterns that are predictive of the target but unrelated to the underlying task. Over-reliance on these shortcuts can lead models to make incorrect predictions under distribution shift, undermining robustness and fairness. While next-token prediction is the canonical pretraining objective for LLMs, it is a difficult setting for analyzing spurious correlations. Here, the label space consists of the full vocabulary, making it difficult to define a clean boundary between meaningful and spurious input features. Consequently, in this paper, we focus on classification-style downstream tasks, where the label space is well-defined and controlled injections of spurious tokens are easier to construct. Adaptation to such tasks is typically done via finetuning. Recently, parameter-efficient finetuning (PEFT) methods like Low-Rank Adaptation (LoRA) have become widely adopted due to their efficiency and scalability. However, real-world datasets are rarely clean. Spurious tokens—such as leftover markup, templated prompts, or systematic metadata patterns—can unintentionally correlate with target labels. Worse yet, malicious actors can intentionally inject such correlations to control the behavior of the finetuned model downstream. If LoRA-finetuned models learn to depend on these shortcuts, it opens the door to test-time manipulation via what we call Seamless Spurious Token Injection (SSTI).
Despite LoRA’s popularity, the interaction between PEFT and spurious correlations remains underexplored. This paper addresses the following two-part research question: (a) Can minimal perturbations—such as a single token of SSTI—suffice to control model behavior? (b) Does LoRA finetuning exacerbate this vulnerability? To study this, we introduce a framework to systematically inject spurious tokens into classification datasets. This enables us to systematically study how models of different sizes and LoRA configurations behave under varying spurious conditions. By isolating key parameters—such as the proportion of affected samples, number of injected tokens, and their placement—we aim to better understand the sensitivity of LoRA-based finetuning to SSTI. We ran comprehensive experiments across three model families (Meta LLaMA-3, Apple OpenELM, and Snowflake Arctic) and four diverse datasets (IMDB, Financial Classification, CommonSenseQA, and Bias in Bios). For clarity and space, the main body of this paper largely focuses on results from the Snowflake Arctic models on the IMDB dataset; however, we observe similar trends across all other models and datasets, with complete results for each experiment provided in the appendix.
We uncover five key findings:
@@
  • Higher LoRA rank surprisingly amplifies susceptibility under light SSTI: When spurious tokens are sparsely injected, the model’s reliance on them increases proportionally with LoRA rank.
@@
Our findings reveal a core weakness in LoRA-based finetuning, raising questions about data quality, model security, and the tradeoff between efficiency and robustness. Especially striking is the observation that a single token and an extremely small number of parameters (e.g., LoRA rank-1) are sufficient to make even a large model (>1B parameters) fully disregard its pretraining knowledge. Alongside this paper, we release a plug-and-play framework for injecting spurious corruptions into Hugging Face datasets, making it to test model robustness as well as facilitate future research on additional corruption strategies (https://github.com/pradyut3501/spurious_corr).
@@ 2 Related Work
The presence of spurious correlations—superficial patterns in the data that models exploit as shortcuts—has been widely documented across both vision and language domains [38]. In computer vision, a canonical example involves classifiers that associate cows with green grass: while models appear to perform well on in-distribution test data, their accuracy collapses on images of cows in atypical contexts, revealing reliance on background texture rather than core object features [11]. In natural language processing (NLP), large language models trained on biased corpora may reinforce social stereotypes, learning shallow associations between demographic terms and harmful concepts rather than robust linguistic generalizations [6]. Recent work has sought to quantify the impact of spurious correlations on model predictions and internal representations [15, 40]. Various testing methodologies have been proposed to detect these correlations, such as evaluating out-of-distribution (OOD) generalization rather than relying solely on in-distribution benchmarks, which may mask shortcut behavior [10, 11]. Other strategies involve curated diagnostic datasets like HANS, designed to expose heuristics in natural language inference models [21]. To address these issues, a wide array of mitigation techniques have been proposed [3, 4, 10, 15, 26, 30, 33, 34, 40]. These fall broadly into two categories: data-centric and model-centric approaches. Data-centric methods include constructing balanced datasets through counterfactual augmentation [40], leveraging human annotation [30], masking previously attended features [4], and reweighting training samples to suppress reliance on spurious signals [10]. Model-centric approaches include deep feature reweighting (DFR)[15], invariant risk minimization (IRM)[3], distributionally robust optimization (DRO)[26], multi-task learning with pretrained models[33], and adversarial training [10]. In particular, DFR, when paired with appropriate architectures and pretraining, has been shown to be highly effective [14]. However, follow-up work has shown that some methods—such as DRO—can fail in the presence of overparameterized models [27], underscoring the need for continued empirical scrutiny.
Parameter Efficient Finetuning
@@
Fine-tuning large language models (LLMs) on downstream tasks can be computationally expensive, especially when dealing with models containing a large number of parameters. To mitigate these costs, a growing body of work has focused on parameter-efficient fine-tuning (PEFT) methods that aim to adapt models with a minimal number of trainable parameters. One of the most prominent approaches is Low-Rank Adaptation (LoRA) [12], which inserts trainable rank-decomposition matrices into the model’s weight updates. LoRA significantly reduces the number of trainable parameters while often achieving performance comparable to, or even surpassing, full fine-tuning. The success of LoRA has led to numerous extensions aimed at further improving efficiency and expressivity. For instance, Nested LoRA (NoRA)[16] builds on singular value decomposition (SVD) to reduce parameter count while preserving information from the base model. Tied-LoRA[25] enhances efficiency through selective training and weight tying. DoRA (Decomposed LoRA) [17] proposes an orthogonal decomposition of the update direction into separate direction and momentum components, further refining the adaptation process. Our work builds on this line of research by examining how PEFT methods—specifically LoRA—respond to training data contaminated with spurious correlations. While prior work has focused on improving adaptation efficiency, we focus on understanding the robustness trade-offs PEFT methods introduce when faced with biased or corrupted training signals.
Malicious Motives
@@
The rise of LLMs has spurred a wave of jailbreak techniques designed to hijack models or bypass their safety measures [5, 7, 8, 18, 24, 28, 29, 32, 35, 36, 39]. Models are vulnerable to various attacks. For example, Wallace et al. show that trigger phrases can control LLM behavior even when not seen during training [35]. AgentPoison compromises RAG-based models by corrupting long-term memory [7], while SequentialBreak hides malicious prompts in long benign sequences to elicit harmful responses [28]. Similarly, a backdoor can be placed in a model during reinforcement learning from human feedback [24]. Shumailov et al. demonstrate that merely changing data order during training—without any injection—can alter a model’s predictions by exploiting stochastic gradient descent [29]. Overall, these techniques are real dangers that have been validated by industry vendors [18].
@@ 3 Method: Seamless Spurious Token Injection (SSTI)
This section introduces the spurious token injection framework that enables our empirical analysis of SSTI (Seamless Spurious Token Injection) introduced in section 1. We begin by formally defining spurious tokens in section 3.1, following which we describe our injection framework in section 3.2. We then detail our experimental setup in section 3.3.
@@ 3.1 A Formalism for Spurious Token Injection
That is, the conditional entropy of the class label given a token in $S$ is substantially lower than for any token outside of $S$. This reflects a strong, potentially unintended association between tokens in $S$ and the target class $y$.
We refer to this as an atomic notion of spuriousness, as it applies at the individual token level, without requiring higher-order interactions or semantic interpretation.
Note. In typical real-world datasets, most tokens are not individually predictive of a label, especially in nontrivial classification tasks. Empirically, this can be validated by computing $H(y\mid t)$ for all tokens $t\in\mathcal{V}$ and observing that the conditional entropy is generally high or near-uniform. See fig. 2 and section A.4 for empirical validation of this. This highlights how atypical it is for a single token to dramatically reduce label uncertainty in well-constructed datasets.
Figure 2: Conditional Entropy for clean IMDB (left, 2 classes) and Common Sense (right, 5 classes) datasets, removing tokens that appear in less than 50 samples. Majority of tokens have a high entropy meaning that their occurrence alone is not enough to predict the prompt class $y$. More examples can be found in fig. 8.
@@
  • Basing our definition on conditional entropy implies a practical way for detecting spurious tokens. This is expanded upon in section 4.6.

### v2

@@ Abstract
Large Language Models (LLMs) are commonly finetuned for a variety of use cases and domains. A common approach is to leverage Low-Rank Adaptation (LoRA)– known to provide strong performance at low resource costs. In this study, we demonstrate that LoRA actually opens the door to short-cut vulnerabilities–and the more resource efficient is the LoRA setup, the more vulnerable will be the finetuned model to aggressive attacks. To measure that vulnerability, we introduce Seamless Spurious Token Injection (SSTI), where we find that LoRA exclusively focuses on even just a single token that is spuriously correlated with downstream labels. In short, injection of that spurious token during finetuning ensure that the model’s prediction at test-time can be manipulated on-demand. We conducted experiments across model families and datasets to evaluate the impact of SSTI during LoRA finetuning while providing possible mitigations. Our experiments conclude that none of the existing checkers and preprocessors can sanitize a dataset raising new concerns for data quality and AI safety.
@@ 1 Introduction
Large language models (LLMs) have achieved impressive performance across a range of natural language processing tasks. However, their generalization can at times be fragile, particularly when training data contains spurious correlations—patterns that are predictive of the target but unrelated to the underlying task. Over-reliance on these shortcuts can lead models to make incorrect predictions under distribution shift, undermining robustness and fairness. While next-token prediction is the canonical pretraining objective for LLMs, it is a difficult setting for analyzing spurious correlations. Here, the label space consists of the full vocabulary, making it difficult to define a clean boundary between meaningful and spurious input features. Consequently, in this paper, we focus on classification-style downstream tasks, where the label space is well-defined and controlled injections of spurious tokens are easier to construct. Adaptation to such tasks is typically done via finetuning. Recently, parameter-efficient finetuning (PEFT) methods like Low-Rank Adaptation (LoRA) have become widely adopted due to their efficiency and scalability. However, real-world datasets are rarely clean. Spurious tokens—such as leftover markup, templated prompts, or systematic metadata patterns—can unintentionally correlate with target labels. Worse yet, malicious actors can intentionally inject such correlations to control the behavior of the finetuned model downstream. If LoRA-finetuned models learn to depend on these shortcuts, it opens the door to test-time manipulation via what we call Seamless Spurious Token Injection (SSTI). Despite LoRA’s popularity, the interaction between PEFT and spurious correlations remains underexplored. This paper addresses the following three-part research question: (a) Can minimal perturbations—such as a single token of SSTI—suffice to control model behavior? (b) Does LoRA finetuning exacerbate this vulnerability? (c) Are existing grammar correction and pre-processing methods effective at mitigating these spurious tokens? To study this, we introduce a framework to systematically inject spurious tokens into classification datasets. This enables us to systematically study how models of different sizes and LoRA configurations behave under varying spurious conditions. By isolating key parameters—such as the proportion of affected samples, number of injected tokens, and their placement—we aim to better understand the sensitivity of LoRA-based finetuning to SSTI. We ran comprehensive experiments across three model families (Meta LLaMA-3, Apple OpenELM, and Snowflake Arctic) and four diverse datasets (IMDB, Financial Classification, CommonSenseQA, and Bias in Bios). Lastly, we leveraged common checkers such as GECTOR (Omelianchuk et al., 2020), T5-GEC (Katinskaia & Yangarber, 2023), and paraphrasing with LLMs in attempts to counter the effects of SSTI.
We uncover some key findings:
@@
  • Semantic integration of spurious elements: Grammar checkers and paraphrasing can misinterpret spurious tokens as valid semantic content, especially entity names and numeral literals, which may unintentionally reinforce misleading correlations.
Our findings reveal a core weakness in LoRA-based finetuning, raising questions about data quality, model security, and the tradeoff between efficiency and robustness. Alongside this paper, we release a plug-and-play framework for injecting spurious corruptions into Hugging Face datasets, making it to test model robustness as well as facilitate future research on additional corruption strategies (https://anonymous.4open.science/r/LLM-research-18B5/README.md).
@@ 2 Related Work
Spurious Correlation: The presence of spurious correlations—superficial patterns in the data that models exploit as shortcuts—has been widely documented across both vision and language domains (Ye et al., 2024). In natural language processing (NLP), large language models trained on biased corpora may reinforce social stereotypes, learning shallow associations between demographic terms and harmful concepts rather than robust linguistic generalizations (Bender et al., 2021). Recent work has sought to quantify the impact of spurious correlations on model predictions and internal representations (Kirichenko et al., 2023; Zhou et al., 2024b; c). Various testing methodologies have been proposed to detect these correlations, such as evaluating out-of-distribution (OOD) generalization rather than relying solely on in-distribution benchmarks, which may mask shortcut behavior (Du et al., 2023; Geirhos et al., 2020). Other strategies involve curated diagnostic datasets like HANS, designed to expose heuristics in natural language inference models (McCoy et al., 2019). To address these issues, a wide array of mitigation techniques have been proposed (Arjovsky et al., 2020; Asgari et al., 2022; Du et al., 2023; Kirichenko et al., 2023; Sagawa et al., 2019; Srivastava et al., 2020; Tu et al., 2020; Varma et al., 2024; Zhou et al., 2024b). These fall broadly into two categories: data-centric and model-centric approaches. Data-centric methods include constructing balanced datasets through counterfactual augmentation (Zhou et al., 2024b), leveraging human annotation (Srivastava et al., 2020), masking previously attended features (Asgari et al., 2022), and reweighting training samples to suppress reliance on spurious signals (Du et al., 2023). Model-centric approaches include deep feature reweighting (DFR)(Kirichenko et al., 2023), invariant risk minimization (IRM)(Arjovsky et al., 2020), distributionally robust optimization (DRO)(Sagawa et al., 2019), multitask learning with pretrained models(Tu et al., 2020), and adversarial training (Du et al., 2023). In particular, DFR, when paired with appropriate architectures and pretraining, has been shown to be highly effective (Izmailov et al., 2022). However, follow-up work has shown that some methods—such as DRO—can fail in the presence of overparameterized models (Sagawa et al., 2020), underscoring the need for continued empirical scrutiny.
@@
Parameter Efficient Finetuning: Fine-tuning large language models (LLMs) on downstream tasks can be computationally expensive. To mitigate these costs, Low-Rank Adaptation (LoRA) (Hu et al., 2021), was proposed which inserts trainable rank-decomposition matrices into the model’s weight updates. LoRA significantly reduces the number of trainable parameters while often achieving performance comparable to full fine-tuning. The success of LoRA has led to numerous extensions. For instance, DoRA (Decomposed LoRA) (Liu et al., 2024a) proposes an orthogonal decomposition of the update direction into separate direction and momentum components We build on this line of research by examining how LoRA responds to training data contaminated with spurious correlations, focusing on understanding the robustness trade-offs LoRA introduces when faced with biased or corrupted training signals.
@@
Malicious Motives: The rise of LLMs has spurred a wave of jailbreak techniques designed to hijack models or bypass their safety measures (Barreno et al., 2006; Chen et al., 2024; Chowdhury et al., 2024; Liu et al., 2024b; Rando & Tramèr, 2024; Saiem et al., 2025; Shumailov et al., 2021; Tian et al., 2024; Wallace et al., 2020; Xu et al., 2024; Zhou et al., 2024a). Models are vulnerable to various attacks. For example, Wallace et al. show that trigger phrases can control LLM behavior even when not seen during training Wallace et al. (2020). AgentPoison compromises RAG-based models by corrupting long-term memory (Chen et al., 2024), while SequentialBreak hides malicious prompts in long benign sequences to elicit harmful responses (Saiem et al., 2025). Similarly, a backdoor can be placed in a model during reinforcement learning from human feedback (Rando & Tramèr, 2024). Shumailov et al. demonstrate that merely changing data order during training—without any injection—can alter a model’s predictions by exploiting stochastic gradient descent Shumailov et al. (2021). Overall, these techniques are real dangers that have been validated by industry vendors (Liu et al., 2024b).
Data Cleaning: Pre-processing and data cleaning are essential steps in most training pipelines. When considering the idea of spurious correlations, we should also pay attention to how they can be affected by data cleaning. If these correlations can be easily removed with existing techniques, then they would be nothing to worry about. We focus predominantly on grammar correction techniques because of the textual nature of our data. Commonly used techniques are GECToR (Omelianchuk et al., 2020) and a finetuned T5 for GEC (Katinskaia & Yangarber, 2023). Recently, LLM paraphrasing has begun to be used as a data augmentation tool and could be applied similarly for preprocessing (Wang et al., 2023).
@@ 3 Method: Seamless Spurious Token Injection (SSTI)
This section introduces the spurious token injection framework that enables our empirical analysis of SSTI (Seamless Spurious Token Injection) introduced in section˜1. We begin by formally defining spurious tokens in section˜3.1, following which we describe our injection framework in section˜3.2. We then detail our experimental setup in section˜3.3.
@@ 3.1 A Formalism for Spurious Token Injection
That is, the conditional entropy of the class label given a token in $S$ is substantially lower than for any token outside of $S$. This reflects a strong, potentially unintended association between tokens in $S$ and the target class $y$. We refer to this as an atomic notion of spuriousness, as it applies at the individual token level, without requiring higher-order interactions or semantic interpretation.
Note. In typical real-world datasets, most tokens are not individually predictive of a label, especially in nontrivial classification tasks. Empirically, this can be validated by computing $H(y\mid t)$ for all tokens $t\in\mathcal{V}$ and observing that the conditional entropy is generally high or near-uniform. See fig.˜2 and section˜A.5 for empirical validation of this. This highlights how atypical it is for a single token to dramatically reduce label uncertainty in well-constructed datasets.
Figure 2: Conditional Entropy for clean IMDB (left, 2 classes) and Common Sense (right, 5 classes) datasets, removing tokens that appear in less than 50 samples. Majority of tokens have a high entropy meaning that their occurrence alone is not enough to predict the prompt class $y$. More examples can be found in fig.˜7.
@@
  • Basing our definition on conditional entropy implies a practical way for detecting spurious tokens. This is expanded upon in section˜5.1.